OpenStudy (anonymous):

breakdown and explanation please solve for x. log25x – log5 = 2 73 + x = 54

OpenStudy (anonymous):

the second equation is not part of it sorry my mistake

OpenStudy (anonymous):

you know that loga - logb = log(a/b) ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

log25x – log5 = 2 log(25x/5) = 2 log(5x) = 2 { when you see log(x) it means log base 10 of x, when you see ln(x) it means log base e of x- if you didn't see 'e' yet , nevermind this } so \[\log_{10}(5x) = 2\] \[5x = 10^2 \]

OpenStudy (anonymous):

the last step explanation is you have a formula: \[\log_{x}a = m \ \ =>\ \ a =m^x \]

OpenStudy (anonymous):

ok so far I am kinda getting this but what formula is that for?

OpenStudy (anonymous):

@SandeepReddy

OpenStudy (anonymous):

we had log(5x) = 2 in the above problem right there \[log_{10}(5x) = 2\] you can see the similarity \[log_{x}(a) = m\]

OpenStudy (anonymous):

oops, i think I confused with the x in formula, did i? { tha't just formula }

OpenStudy (anonymous):

im confused now haha

OpenStudy (anonymous):

http://www.mathwords.com/l/logarithm_rules.htm

OpenStudy (anonymous):

you can get it from this way 2 = log 100 because 10^2 =100 hence rewrite as \(log(5x) = log (100)\) hence 5x =100, x=20